# Deformation and breakup of the liquid ligament with various disturbances on the interface in shear flow

### v1

@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
Figure 10 shows the position $y/d$ where the interface at each $\Delta U/U_{\mathrm{ref}}$ is the highest at the time just before the splitting of the liquid ligament. For $\Delta U/U_{\mathrm{ref}}=5.9$ and 11.9, the interface around $z/\lambda=0$, 0.5, and 1.0 bulges toward the gas side. This result is due to a nonlinear effect, and the turbulence with twice the wavenumber of the initial disturbance occurs at the interface. For $\Delta U/U_{\mathrm{ref}}=17.8$, in addition to $z/\lambda=0$, 0.5, and 1.0, the interface also expands to the gas side around $z/\lambda=0.25$ and 0.75. Compared with the results of $\Delta U/U_{\mathrm{ref}}=5.9$ and 11.9, higher wavenumber turbulence occurs at the interface. This high wavenumber turbulence grows, and the liquid ligament in the vicinity indicated by the arrows in Fig. 9(c) are constricted and split, resulting in an increase in the number of breakup droplets for $\Delta U/U_{\mathrm{ref}}=17.8$.
@@
Figure 11 shows the diameters $D$ of the droplets that broke up for $\Delta U/U_{\rm ref}=5.9$, 11.9, and 17.8. The coordinate $z/\lambda$ represents the location of the droplet. In this study, similar to the existing study (Yanaoka and Nakayama, 2022), the droplet diameter is defined as the equivalent diameter obtained from the droplet volume. The solid and dashed lines in the figure are the theoretical values for an inviscid liquid column (Rayleigh, 1878) and for a viscous liquid column (Weber, 1931), respectively. Weber’s theoretical value is expressed as follows:
@@
The droplet size in this calculation is smaller than the two theoretical values (Rayleigh, 1878; Weber, 1931). The theoretical values are based on linear theory and do not consider the nonlinear effects of flow. In this calculation, as shown in Fig. 10, turbulence with higher wavenumber components than the initial disturbance occurs at the interface due to the nonlinear effect. It is considered that as the liquid ligament with a short wavelength breaks up due to this high wavenumber turbulence, the droplet diameter is smaller than the theoretical value (Rayleigh, 1878; Weber, 1931). As $\Delta U/U_{\mathrm{ref}}$ increases, the droplet diameters at $z/\lambda=0$ and 0.5 decrease and increase, respectively. The ratio of the droplet diameters of $z/\lambda=0.5$ and $z/\lambda=0$, $D_{z/\lambda=0.5}$/$D_{z/\lambda=0}$, is 0.37, 0.96, and 0.98 for $\Delta U/U_{\mathrm{ref}}=5.9$, 11.9 and 17.8, respectively. It can be seen that the droplet diameter becomes uniform as $\Delta U/U_{\mathrm{ref}}$ increases. In addition, for $\Delta U/U_{\mathrm{ref}}=17.8$, the total surface area of the liquid increases because the number of breakup droplets increases. Therefore, it can be said that increasing $\Delta U/U_{\mathrm{ref}}$ improved the atomization quality of the liquid ligament.
@@
In actual flow fields, gas and liquid flows are turbulent, and therefore, turbulences with various wavenumber components occur at the gas–liquid interface. In this analysis, we investigate the effect of such disturbances on the deformation and splitting of a liquid ligament. First, for the velocity difference $\Delta U/U_{\mathrm{ref}}=17.8$, where the influence of the shear velocity is the most intensive, we vary the dimensionless wavenumber $ka$ as 0.3, 0.5, 0.7, and 0.9. We investigate the effect of the wavenumber of the initial disturbance at the interface on the deformation and splitting of the liquid ligament.
Figures12 to 14 show the time variation of the liquid ligament interface at $ka=0.3$, 0.5, and 0.9 for $\Delta U/U_{\mathrm{ref}}=17.8$. The result at $ka=0.7$ was shown in Fig. 8. At $T=0.6$, the liquid ligaments for all $ka$ are deformed into films by the shear flow. As $ka$ increases, the portions of $z/\lambda=0$ and 0.5 in the liquid ligament bulge. At $T=3.0$ and $T=2.3$ for $ka=0.3$ and 0.5, respectively, turbulence with a higher wavenumber than the initial disturbance occurs at the interface of $z/\lambda=0.1-0.4$ and $0.6-0.9$. At $T=4.3$, 3.0, 2.7, and 2.1 for $ka=0.3$, 0.5, 0.7, and 0.9, respectively, the liquid ligament splits, and the breakup time of the liquid ligament shortens as $ka$ increases.
Figure 15 shows the time variation of the $z$-axial velocity $w_{\mathrm{av}}$ averaged in the cross-section of the liquid ligament at $z/\lambda=0.6$ and 0.9. At $z/\lambda=0.6$ and 0.9, the liquids flow in the negative and positive directions of the $z$-axis, respectively. The absolute value of $w_{\mathrm{av}}$ increases over time for all $ka$ at the two cross-sections. In addition, the magnitude of $w_{\mathrm{av}}$ increases as $ka$ increases. It can be seen from this that the moving liquid velocity along the central axis of the liquid ligament differs depending on $ka$, and the moving speed increases as $ka$ increases.
@@
We investigate the change in the flow field at $T=1.0$, where the difference in the interface shape due to $ka$ becomes large. Figure 16 shows the interface and three-dimensional velocity vectors for each $ka$. The velocity vector shows the distribution of the cross-section across the liquid ligament. As $ka$ increases, the bulging parts at $z/\lambda=0$ and 0.5 in the liquid ligament grow, and the bulge shape becomes spherical. In addition, the liquid ligament near $z/\lambda=0.25$ and 0.75 is more constricted. At $ka=0.3$ and 0.5 on the low wavenumber side, a three-dimensional vortex is formed at the constriction. Near the bulge at $z/\lambda=0.5$, a flow toward $z/\lambda=0.5$ occurs. On the other hand, at $ka=0.7$ and 0.9 on the high wavenumber side, a three-dimensional vortex is formed near the bulge at $z/\lambda=0$, 0.5, and 1.0. Due to these vortices, the liquid inside the liquid ligament flows from the constricted part to the bulging part, and the bulging part develops.
@@
For all $ka$, the cross-sectional area of the liquid ligament at $z/\lambda=0.75$ decreases with time. As $ka$ increases, the cross-sectional area decreases rapidly. As shown in Fig. 15, with an increase in $ka$, the liquid movement along the central axis of the liquid ligament is promoted. As shown in Fig. 16, on the high wavenumber side, the vortices that enhance the development of the bulge are formed near the interface. As an increase in $ka$ shortens the wavelength, the liquid inside the ligament shortly moves from the constriction to the bulge. Therefore, as $ka$ increases, the cross-sectional area of the liquid ligament at $z/\lambda=0.75$ becomes small in a short time, and the breakup time of the liquid ligament becomes short.
@@
As mentioned, a liquid ligament breaks up faster in an initial disturbance with a high wavenumber than a low wavenumber. On the other hand, for low wavenumbers, as shown in Figs. 12(f) and 13(d), new turbulence with a higher wavenumber than the initial disturbance occurs at the interface. To quantitatively investigate the difference in the interface shape due to $ka$, Fig. 18 shows the time variation in the interface for each $ka$. The distribution is the interface at the position $y/d$ where the interface is the highest in the $x$$y$ cross-section. At all times for $ka=0.9$, the interface shape is similar, and the initial disturbance grows. At $T=1.0$ and 1.5 for $ka=0.5$ and 0.7, respectively, the parts at $z/\lambda=0$ and 0.5 of the liquid ligament bulge toward the gas side. That is, turbulence with twice the wavenumber of the initial disturbance occurs at the interface. Moreover, the interface expands with time toward the gas side even around $z/\lambda=0.25$ and 0.75, and at $T=2.3$, turbulence with four times the initial disturbance occurs at the interface. At $T=1.0$ for $ka=0.3$, the parts at $z/\lambda=0$ and 0.5 of the liquid ligament bulge toward the gas side, and turbulence with twice the initial disturbance occurs at the interface. At $T=2.4$, the interface bulges slightly toward the gas side near $z/\lambda=0.12$, 0.35, 0.65, and 0.88. The raised interface develops over time. At $T=2.8$, near $z/\lambda=0.15$, 0.35, 0.65, and 0.85, the interface rises more toward the gas side, and turbulence with a wavenumber six times that of the initial disturbance occurs at the interface. For $ka=0.3$, the uplift of the interface toward the gas side is the largest, and the nonlinear effect appears most strongly.
It was found from the above results that the liquid ligament deforms more complicatedly for the initial disturbance with a low wavenumber than a high wavenumber due to a nonlinear effect. Therefore, we considered that the deformation of the liquid ligament would be accelerated by adding the high wavenumber disturbance to the low wavenumber disturbance. For $\Delta U/U_{\mathrm{ref}}=17.8$ and $ka=0.3$, we investigate the effects of low and high wavenumber disturbances on the deformation and splitting of the ligament.
@@
Figures 19 and 20 show the time evolution of the liquid ligament interface for $\eta_{r}=0.9$ at $k_{r}=2.0$ and 4.0, respectively. At $T=0.3$ for $k_{r}=2.0$, the liquid ligament is deformed into a film by the influence of the shear flow. At $T=1.0$, perforation occurs at $z/\lambda=0$ in the liquid ligament, but at $T=1.6$, the liquid is combined. At $T=2.6$, turbulence with a higher wavenumber than the initial disturbance occurs, similar to the case of $\eta_{r}=0$ and $k_{r}=0$ shown in Fig. 12. When time passes until $T=4.6$, the splitting of the liquid ligament is confirmed, and large and small liquid droplets are formed near $z/\lambda=0$, 0.5 and $z/\lambda=0.2$, 0.8, respectively. For $k_{r}=4.0$, the liquid ligament is deformed into a film at $T=0.3$, similar to the result for $k_{r}=2.0$. The parts at $z/\lambda=0$, 0.25, 0.5, and 0.75 of the liquid ligament bulge due to the initial disturbance and swell further with time. At $T=1.0$, perforation occurs at $z/\lambda=0$ in the liquid ligament. At $T=2.0$, the part at $z/\lambda=0$ of the liquid ligament is split by the shear flow. The breakup droplets flow in the negative direction of the $x$-axis due to the airflow. At $T=2.0$, turbulence with a higher wavenumber than the initial disturbance occurs at $z/\lambda=0.1-0.4$ and $0.6-0.9$. The bulge formed by this high wavenumber turbulence develops with time, and at $T=3.0$, the liquid ligament breaks up. The ligament breakup time for $\eta_{r}=0.9$, $k_{r}=4.0$ becomes shorter than those for $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$. In addition, the number of breakup droplets increases.
@@
Figure 21 shows the time variation of the interface for each $\eta_{r}$ and $k_{r}$. The result for $\eta_{r}=0$ and $k_{r}=0$ were shown in Fig. 18(a). The distribution is the interface at the position $y/d$ where the interface is the highest in the $x$$y$ cross-section. At $T=2.0$ for $\eta_{r}=0.9$ and $k_{r}=4.0$, the position of the droplet interface near $z/\lambda=0$ is higher than that of the liquid ligament interface near $z/\lambda=0$. Thus, except for this droplet, we show the position $y/d$ where the liquid ligament interface is the highest. For $k_{r}=2.0$, the initial disturbance grows at $T=1.0$, but no new turbulence with high wavenumber occurs. At $T=2.0$, the interface near $z/\lambda=0.18$, 0.38, 0.68, and 0.82 slightly rises toward the gas side. At $T=2.6$, this bulge develops, and turbulence with a wavenumber six times that of the initial disturbance is generated at the interface. This trend is similar to the results for $\eta_{r}=0$ and $k_{r}=0$. Even for $k_{r}=4.0$, the initial disturbance grows at $T=1.0$, but no new turbulence with high wavenumber occurs. At $T=1.6$, the interface near $z/\lambda=0$, 0.15, 0.5, 0.85, and 1.0 rises toward the gas side. It is also found that the interface near $z/\lambda=0.38$ and 0.68 slightly bulges toward the gas side. At $T=2.0$, two portions of the interface at $z/\lambda=0.1-0.4$ and $0.6-0.9$ rise toward the gas side, and new turbulence with a higher wavenumber than the initial disturbance appears. For $\eta_{r}=0.9$ and $k_{r}=4.0$, higher wavenumber turbulence than the initial disturbance occurs on the interface at earlier times than the other conditions, and the bulge of the interface toward the gas side also increases. Therefore, we can see that the nonlinear effect appears stronger than the results for $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$. As shown in Fig. 20, depending on the strength of the nonlinear effect, the deformation process of the liquid ligament, the number of breakup droplets, and breakup time are different from the results for $\eta_{r}=0$, $k_{r}=0$ and for $\eta_{r}=0.9$, $k_{r}=2.0$. In the following, we investigate in detail the difference in the deformation of liquid ligaments between $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=4.0$.
@@
Figure 22 shows the interface, velocity vectors, $y$-direction vorticity, and streamlines in the $x$$z$ cross-section at $y/d=\pm{0.04}$ for $\eta_{r}=0.9$ and $k_{r}=4.0$ at $T=0.3$. At $x/d>0$ near $z/\lambda=0.5$, the airflow flows in the positive direction of the $x$-axis. At $x/d<0$ near $z/\lambda=0$ and 1.0, the airflow flows in the negative direction of the $x$-axis. From the streamlines, we can also see that at $y/d=0.04$, vortices due to airflow are formed near $(x/d,z/\lambda)=(1.3,0.1)$ and $(1.3,0.9)$. On the other hand, at $y/d=-0.04$, vortices due to airflow are formed near $(x/d,z/\lambda)=(-1.3,0.35)$ and $(-1.3,0.65)$. These vortices form positive and negative vorticities in the regions around $(x/d,z/\lambda)=(1.5,0.3)$, $(-1.5,0.2)$ and $(x/d,z/\lambda)=(1.5,0.7)$, $(-1.5,0.8)$, respectively. The parts around $(x/d,z/\lambda)=(1.25,0.25)$, $(-1.25,0.25)$, $(1.25,0.75)$, and $(-1.25,0.75)$ of the liquid ligament rise. It is considered that the vortices move these bulge parts along the central axis of the liquid ligament.
@@
Figure 23 shows the time evolution of the interface and velocity vector for $\eta_{r}=0.9$ and $k_{r}=4.0$. Part A, shown in the left figure, is the bulge existing at $x/d>0$ around $z/\lambda=0.25$ and 0.75 in Fig. 22. Part B, shown in the right figure, is the bulge existing at $x/d<0$ around $z/\lambda=0.25$ and 0.75 in Fig. 22. Velocity vectors in the $y$$z$ cross-section where the bulge parts A and B are present are shown. At all times, the center of the vortex around the liquid ligament exists inside the ligament. In addition, all bulges develop with time. Around the interface of the bulge part A at $T=1.0$, the vortex flows the airflow toward $z/\lambda=0.5$. On the other hand, near the interface of the bulge part B, the vortex flows the airflow toward $z/\lambda=0$, 1.0. Therefore, at $T=1.4$, the bulges A and B move to $z/\lambda=0.5$ and $z/\lambda=0$, 1.0, respectively. At $T=1.6$, unlike $T=1.0$, near the interface of the bulge A, the airflow flows toward $z/\lambda=0$, 1.0 due to the vortex. Near the interface of the bulge B, the vortex causes the airflow to flow toward $z/\lambda=0.5$. Therefore, at $T=1.8$, the bulges A and B move to $z/\lambda=0$, 1.0 and $z/\lambda=0.5$, respectively. In this way, the bulging part of the liquid due to the initial disturbance is moved by the vortex with time along the central axis of the liquid ligament.
For $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=4.0$, Fig. 24 shows the time variation of the cross-sectional area $S$. In Fig. 24(a), the cross-section areas at $z/\lambda=0$ and 0.5 increase with time. On the other hand, the areas at $z/\lambda=0.25$ and 0.75 decrease with time. As shown in Fig. 22, at $T=0.3$ for $\eta_{r}=0.9$ and $k_{r}=4.0$, the parts at $z/\lambda=0.25$ and 0.75 swell due to the initial disturbance. Therefore, in Fig. 24(b), the cross sections near $z/\lambda=0.25$ and 0.75 are larger than those near $z/\lambda=0.1$, 0.4, 0.6, and 0.9. The vortices move the bulges at $z/\lambda=0.1-0.4$ and $0.6-0.9$ along the central axis of the liquid ligament. Thus, at $T=1.0$, the cross sections at $z/\lambda=0.15$, 0.35, 0.65, and 0.85 are larger than those at $z/\lambda=0.25$ and 0.75. With time, the bulge deforms into a spherical shape due to surface tension, and the liquid ligament around the bulge becomes constricted. At $T=1.8$, the cross sections at $z/\lambda=0.25$ and 0.75 for $\eta_{r}=0.9$ and $k_{r}=4.0$ are smaller than those at $z/\lambda=0.25$ and 0.75 for $\eta_{r}=0$ and $k_{r}=0$. It can be seen from this that the deformation of the liquid ligament is enhanced for $\eta_{r}=0.9$ and $k_{r}=4.0$.

### v2

@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
Figure 10 shows the position $y/d$ where the interface at each $\Delta U/U_{\mathrm{ref}}$ is the highest at the time just before the splitting of the liquid ligament. For $\Delta U/U_{\mathrm{ref}}=5.9$ and 11.9, the interface around $z/\lambda=0$, 0.5, and 1.0 bulges toward the gas side. This result is owing to a nonlinear effect, and the turbulence with twice the wavenumber of the initial disturbance occurs at the interface. For $\Delta U/U_{\mathrm{ref}}=17.8$, in addition to $z/\lambda=0$, 0.5, and 1.0, the interface also expands to the gas side around $z/\lambda=0.25$ and 0.75. Compared with the results of $\Delta U/U_{\mathrm{ref}}=5.9$ and 11.9, higher wavenumber turbulence occurs at the interface. This high wavenumber turbulence grows, and the liquid ligament in the vicinity indicated by the arrows in Fig. 9(c) are constricted and split, resulting in an increase in the number of breakup droplets for $\Delta U/U_{\mathrm{ref}}=17.8$.
@@
Figure 11 shows the diameters $D$ of the droplets that broke up for $\Delta U/U_{\rm ref}=5.9$, 11.9, and 17.8. The coordinate $z/\lambda$ represents the location of the droplet. In this study, as in the existing report Yanaoka and Nakayama (2022), the droplet diameter is defined as the equivalent diameter obtained from the droplet volume. The solid and dashed lines in the figure are the theoretical values for an inviscid liquid column (Rayleigh, 1878) and for a viscous liquid column (Weber, 1931), respectively. Weber’s theoretical value is expressed as follows:
@@
The droplet size in this calculation is smaller than the two theoretical values (Rayleigh, 1878; Weber, 1931). The theoretical values are based on linear theory and do not consider the nonlinear effects of flow. In this calculation, as shown in Fig. 10, turbulence with higher wavenumber components than the initial disturbance occurs at the interface owing to the nonlinear effect. It is considered that as the liquid ligament with a short wavelength breaks up owing to this high wavenumber turbulence, the droplet diameter is smaller than the theoretical value (Rayleigh, 1878; Weber, 1931). When $\Delta U/U_{\mathrm{ref}}$ increases, the droplet diameters at $z/\lambda=0$ and 0.5 decrease and increase, respectively. The ratio of the droplet diameters of $z/\lambda=0.5$ and $z/\lambda=0$, $D_{z/\lambda=0.5}$/$D_{z/\lambda=0}$, is 0.37, 0.96, and 0.98 for $\Delta U/U_{\mathrm{ref}}=5.9$, 11.9 and 17.8, respectively. It can be seen that the droplet diameter becomes uniform as $\Delta U/U_{\mathrm{ref}}$ increases. Additionally, for $\Delta U/U_{\mathrm{ref}}=17.8$, the number of breakup droplets increases; hence, the total surface area of the liquid increases. Therefore, it can be said that increasing $\Delta U/U_{\mathrm{ref}}$ improved the atomization quality of the liquid ligament.
@@
In actual flow fields, gas and liquid flows are turbulent; hence, turbulences with various wavenumber components occur at the gas–liquid interface. In this analysis, we investigate the effect of such disturbances on the deformation and splitting of a liquid ligament. First, for the velocity difference $\Delta U/U_{\mathrm{ref}}=17.8$, where the influence of the shear velocity is the most intensive, we vary the dimensionless wavenumber $ka$ as 0.3, 0.5, 0.7, and 0.9. We investigate the effect of the wavenumber of the initial disturbance at the interface on the deformation and splitting of the liquid ligament.
Figures12 to 14 show the time variation of the liquid ligament interface at $ka=0.3$, 0.5, and 0.9 for $\Delta U/U_{\mathrm{ref}}=17.8$. The result at $ka=0.7$ was shown in Fig. 8. At $T=0.6$, the liquid ligaments for all $ka$ are deformed into films by the shear flow. With an increase in $ka$, the portions of $z/\lambda=0$ and 0.5 in the liquid ligament bulge. At $T=3.0$ and $T=2.3$ for $ka=0.3$ and 0.5, respectively, turbulence with a higher wavenumber than the initial disturbance occurs at the interface of $z/\lambda=0.1-0.4$ and $0.6-0.9$. At $T=4.3$, 3.0, 2.7, and 2.1 for $ka=0.3$, 0.5, 0.7, and 0.9, respectively, the liquid ligament splits, and the breakup time of the liquid ligament shortens as $ka$ increases.
Figure 15 shows the time variation of the $z$-axial velocity $w_{\mathrm{av}}$ averaged in the cross-section of the liquid ligament at $z/\lambda=0.6$ and 0.9. At $z/\lambda=0.6$ and 0.9, the liquids flow in the negative and positive directions of the $z$-axis, respectively. The absolute value of $w_{\mathrm{av}}$ increases over time for all $ka$ at the two cross-sections. Additionally, the magnitude of $w_{\mathrm{av}}$ increases as $ka$ increases. It can be seen from this that the moving liquid velocity along the central axis of the liquid ligament differs depending on $ka$, and the moving speed increases as $ka$ increases.
@@
We investigate the change in the flow field at $T=1.0$, where the difference in the interface shape owing to $ka$ becomes large. Figure 16 shows the interface and three-dimensional velocity vectors for each $ka$. The velocity vector shows the distribution of the cross-section across the liquid ligament. As $ka$ increases, the bulging parts at $z/\lambda=0$ and 0.5 in the liquid ligament grow, and the bulge shape becomes spherical. Additionally, the liquid ligament near $z/\lambda=0.25$ and 0.75 is more constricted. At $ka=0.3$ and 0.5 on the low wavenumber side, a three-dimensional vortex is formed at the constriction. Near the bulge at $z/\lambda=0.5$, a flow toward $z/\lambda=0.5$ occurs. Conversely, at $ka=0.7$ and 0.9 on the high wavenumber side, a three-dimensional vortex is formed near the bulge at $z/\lambda=0$, 0.5, and 1.0. Because of these vortices, the liquid inside the liquid ligament flows from the constricted part to the bulging part, and the bulging part develops.
@@
For all $ka$, the cross-sectional area of the liquid ligament at $z/\lambda=0.75$ decreases with time. As $ka$ increases, the cross-sectional area decreases rapidly. As shown in Fig. 15, with an increase in $ka$, the liquid movement along the central axis of the liquid ligament is promoted. As shown in Fig. 16, on the high wavenumber side, the vortices that enhance the development of the bulge are formed near the interface. An increase in $ka$ shortens the wavelength; hence, the liquid inside the ligament shortly moves from the constriction to the bulge. Therefore, as $ka$ increases, the cross-sectional area of the liquid ligament at $z/\lambda=0.75$ becomes small in a short time, and the breakup time of the liquid ligament becomes short.
@@
As mentioned, a liquid ligament breaks up faster in an initial disturbance with a high wavenumber than a low wavenumber. Conversely, for low wavenumbers, as shown in Figs. 12(f) and 13(d), new turbulence with a higher wavenumber than the initial disturbance occurs at the interface. To quantitatively investigate the difference in the interface shape owing to $ka$, Fig. 18 shows the time variation in the interface for each $ka$. The distribution is the interface at the position $y/d$ where the interface is the highest in the $x$$y$ cross-section. At all times for $ka=0.9$, the interface shape is similar, and the initial disturbance grows. At $T=1.0$ and 1.5 for $ka=0.5$ and 0.7, respectively, the parts at $z/\lambda=0$ and 0.5 of the liquid ligament bulge toward the gas side. That is, turbulence with twice the wavenumber of the initial disturbance occurs at the interface. Moreover, the interface expands with time toward the gas side even around $z/\lambda=0.25$ and 0.75, and at $T=2.3$, turbulence with four times the initial disturbance occurs at the interface. At $T=1.0$ for $ka=0.3$, the parts at $z/\lambda=0$ and 0.5 of the liquid ligament bulge toward the gas side, and turbulence with twice the initial disturbance occurs at the interface. At $T=2.4$, the interface bulges slightly toward the gas side near $z/\lambda=0.12$, 0.35, 0.65, and 0.88. The raised interface develops over time. At $T=2.8$, near $z/\lambda=0.15$, 0.35, 0.65, and 0.85, the interface rises more toward the gas side, and turbulence with a wavenumber six times that of the initial disturbance occurs at the interface. For $ka=0.3$, the uplift of the interface toward the gas side is the largest, and the nonlinear effect appears most strongly.
It was found from the above results that the liquid ligament deforms more complicatedly for the initial disturbance with a low wavenumber than a high wavenumber owing to a nonlinear effect. Therefore, we considered that the deformation of the liquid ligament would be accelerated by adding the high wavenumber disturbance to the low wavenumber disturbance. For $\Delta U/U_{\mathrm{ref}}=17.8$ and $ka=0.3$, we investigate the effects of low and high wavenumber disturbances on the deformation and splitting of the ligament.
@@
Figures 19 and 20 show the time evolution of the liquid ligament interface for $\eta_{r}=0.9$ at $k_{r}=2.0$ and 4.0, respectively. At $T=0.3$ for $k_{r}=2.0$, the liquid ligament is deformed into a film by the influence of the shear flow. At $T=1.0$, perforation occurs at $z/\lambda=0$ in the liquid ligament, but at $T=1.6$, the liquid is combined. At $T=2.6$, turbulence with a higher wavenumber than the initial disturbance occurs, similarly to the case of $\eta_{r}=0$ and $k_{r}=0$ shown in Fig. 12. When time passes until $T=4.6$, the splitting of the liquid ligament is confirmed, and large and small liquid droplets are formed near $z/\lambda=0$, 0.5 and $z/\lambda=0.2$, 0.8, respectively. For $k_{r}=4.0$, the liquid ligament is deformed into a film at $T=0.3$, similarly to the result for $k_{r}=2.0$. The parts at $z/\lambda=0$, 0.25, 0.5, and 0.75 of the liquid ligament bulge under the initial disturbance and swell further with time. At $T=1.0$, perforation occurs at $z/\lambda=0$ in the liquid ligament. At $T=2.0$, the part at $z/\lambda=0$ of the liquid ligament is split by the shear flow. The breakup droplets flow in the negative direction of the $x$-axis owing to the airflow. At $T=2.0$, turbulence with a higher wavenumber than the initial disturbance occurs at $z/\lambda=0.1-0.4$ and $0.6-0.9$. The bulge formed by this high wavenumber turbulence develops with time, and at $T=3.0$, the liquid ligament breaks up. The ligament breakup time for $\eta_{r}=0.9$, $k_{r}=4.0$ becomes shorter than those for $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$. Additionally, the number of breakup droplets increases.
@@
Figure 21 shows the time variation of the interface for each $\eta_{r}$ and $k_{r}$. The result for $\eta_{r}=0$ and $k_{r}=0$ were shown in Fig. 18(a). The distribution is the interface at the position $y/d$ where the interface is the highest in the $x$$y$ cross-section. At $T=2.0$ for $\eta_{r}=0.9$ and $k_{r}=4.0$, the position of the droplet interface near $z/\lambda=0$ is higher than that of the liquid ligament interface near $z/\lambda=0$. Hence, except for this droplet, we show the position $y/d$ where the liquid ligament interface is the highest. For $k_{r}=2.0$, the initial disturbance grows at $T=1.0$, but no new turbulence with high wavenumber occurs. At $T=2.0$, the interface near $z/\lambda=0.18$, 0.38, 0.68, and 0.82 slightly rises toward the gas side. At $T=2.6$, this bulge develops, and turbulence with a wavenumber six times that of the initial disturbance is generated at the interface. This trend is similar to the results for $\eta_{r}=0$ and $k_{r}=0$. Even for $k_{r}=4.0$, the initial disturbance grows at $T=1.0$, but no new turbulence with high wavenumber occurs. At $T=1.6$, the interface near $z/\lambda=0$, 0.15, 0.5, 0.85, and 1.0 rises toward the gas side. It is also found that the interface near $z/\lambda=0.38$ and 0.68 slightly bulges toward the gas side. At $T=2.0$, two portions of the interface at $z/\lambda=0.1-0.4$ and $0.6-0.9$ rise toward the gas side, and new turbulence with a higher wavenumber than the initial disturbance appears. For $\eta_{r}=0.9$ and $k_{r}=4.0$, higher wavenumber turbulence than the initial disturbance occurs on the interface at earlier times than the other conditions, and the bulge of the interface toward the gas side also increases. Therefore, we can see that the nonlinear effect appears stronger than the results for $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=2.0$. As shown in Fig. 20, depending on the strength of the nonlinear effect, the deformation process of the liquid ligament, the number of breakup droplets, and breakup time are different from the results for $\eta_{r}=0$, $k_{r}=0$ and for $\eta_{r}=0.9$, $k_{r}=2.0$. In the following, we investigate in detail the difference in the deformation of liquid ligaments between $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=4.0$.
@@
Figure 22 shows the interface, velocity vectors, $y$-direction vorticity, and streamlines in the $x$$z$ cross-section at $y/d=\pm{0.04}$ for $\eta_{r}=0.9$ and $k_{r}=4.0$ at $T=0.3$. At $x/d>0$ near $z/\lambda=0.5$, the airflow flows in the positive direction of the $x$-axis. At $x/d<0$ near $z/\lambda=0$ and 1.0, the airflow flows in the negative direction of the $x$-axis. From the streamlines, we can also see that at $y/d=0.04$, vortices owing to airflow are formed near $(x/d,z/\lambda)=(1.3,0.1)$ and $(1.3,0.9)$. At $y/d=-0.04$, vortices owing to airflow are formed near $(x/d,z/\lambda)=(-1.3,0.35)$ and $(-1.3,0.65)$. These vortices form positive and negative vorticities in the regions around $(x/d,z/\lambda)=(1.5,0.3)$, $(-1.5,0.2)$ and $(x/d,z/\lambda)=(1.5,0.7)$, $(-1.5,0.8)$, respectively. The parts around $(x/d,z/\lambda)=(1.25,0.25)$, $(-1.25,0.25)$, $(1.25,0.75)$, and $(-1.25,0.75)$ of the liquid ligament rise. It is considered that the vortices move these bulge parts along the central axis of the liquid ligament.
@@
Figure 23 shows the time evolution of the interface and velocity vector for $\eta_{r}=0.9$ and $k_{r}=4.0$. Part A, shown in the left figure, is the bulge existing at $x/d>0$ around $z/\lambda=0.25$ and 0.75 in Fig. 22. Part B, shown in the right figure, is the bulge existing at $x/d<0$ around $z/\lambda=0.25$ and 0.75 in Fig. 22. Velocity vectors in the $y$$z$ cross-section where the bulge parts A and B are present are shown. At all times, the center of the vortex around the liquid ligament exists inside the ligament. Additionally, all bulges develop with time. Around the interface of the bulge part A at $T=1.0$, the vortex flows the airflow toward $z/\lambda=0.5$. Near the interface of the bulge part B, the vortex flows the airflow toward $z/\lambda=0$, 1.0. Thus, at $T=1.4$, the bulges A and B move to $z/\lambda=0.5$ and $z/\lambda=0$, 1.0, respectively. At $T=1.6$, unlike $T=1.0$, near the interface of the bulge A, the airflow flows toward $z/\lambda=0$, 1.0 owing to the vortex. Near the interface of the bulge B, the vortex causes the airflow to flow toward $z/\lambda=0.5$. Thus, at $T=1.8$, the bulges A and B move to $z/\lambda=0$, 1.0 and $z/\lambda=0.5$, respectively. In this way, the bulging part of the liquid owing to the initial disturbance is moved by the vortex with time along the central axis of the liquid ligament.
For $\eta_{r}=0$, $k_{r}=0$ and $\eta_{r}=0.9$, $k_{r}=4.0$, Fig. 24 shows the time variation of the cross-sectional area $S$. In Fig. 24(a), the cross-section areas at $z/\lambda=0$ and 0.5 increase with time. Conversely, the areas at $z/\lambda=0.25$ and 0.75 decrease with time. As shown in Fig. 22, at $T=0.3$ for $\eta_{r}=0.9$ and $k_{r}=4.0$, the parts at $z/\lambda=0.25$ and 0.75 swell owing to the initial disturbance. Hence, in Fig. 24(b), the cross sections near $z/\lambda=0.25$ and 0.75 are larger than those near $z/\lambda=0.1$, 0.4, 0.6, and 0.9. The vortices move the bulges at $z/\lambda=0.1-0.4$ and $0.6-0.9$ along the central axis of the liquid ligament. Thus, at $T=1.0$, the cross sections at $z/\lambda=0.15$, 0.35, 0.65, and 0.85 are larger than those at $z/\lambda=0.25$ and 0.75. With time, the bulge deforms into a spherical shape owing to surface tension, and the liquid ligament around the bulge becomes constricted. At $T=1.8$, the cross sections at $z/\lambda=0.25$ and 0.75 for $\eta_{r}=0.9$ and $k_{r}=4.0$ are smaller than those at $z/\lambda=0.25$ and 0.75 for $\eta_{r}=0$ and $k_{r}=0$. It can be seen from this that the deformation of the liquid ligament is enhanced for $\eta_{r}=0.9$ and $k_{r}=4.0$.